solve the differential equation dy/dx + (x/1-x^2)y=xy^1/2 , y(0)=1


1
Expert's answer
2021-04-15T07:19:26-0400

Given question


y'{1−x2x​ } = xy1/2


"\implies" "dy\over \sqrt{y}" =(1-x2)dx


"\implies" "\intop" "dy\over \sqrt{2}" ="\intop" (1-x2)dx


"\implies" 2y1/2 = x - "x^3\over 3" + C

Given y(0) = 1

"\implies" at x= 0 , y = 1


"\implies" 2 = 0 - 0 + C

"\rightarrow" C=2


Answer

"\boxed {2y^{1\over 2 }= x - {x^3\over 3} +2 }"